OpenStudy (anonymous):

Marcus cut pieces of ribbon 4 inches long to decorate name tags for the teacher conference. Which equation can be used to find t, the number of name tags Marcus can decorate with the ribbon from a roll 52 inches long? Marcus cut pieces of ribbon 4 inches long to decorate name tags for the teacher conference. Which equation can be used to find t, the number of name tags Marcus can decorate with the ribbon from a roll 52 inches long? @Mathematics

OpenStudy (anonymous):

how to start this one @ wlbrt

OpenStudy (anonymous):

t/4=52=52

OpenStudy (anonymous):

4t=52

OpenStudy (anonymous):

here as it has been given t number of teachers are there so

OpenStudy (anonymous):

52/4=t t=13

OpenStudy (anonymous):

for 1 teacher ribbon is 4 inches so for t teachers it would be equal to = ??????????

OpenStudy (anonymous):

I'm 5th grader but I think is t/4=52=52

OpenStudy (anonymous):

or 4t=52

OpenStudy (anonymous):

for 1 person = 4 inches thus, for t persons = 4t inches right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Some of these kids are in grade 5???

OpenStudy (anonymous):

so, 4t = total length of ribbon

OpenStudy (anonymous):

now find the value of t

OpenStudy (anonymous):

4t=52

OpenStudy (anonymous):

yes so t = ??

OpenStudy (anonymous):

Sheg you can be a teacher!!!

OpenStudy (anonymous):

54

OpenStudy (anonymous):

why rld ??

OpenStudy (anonymous):

how it will be 54

OpenStudy (anonymous):

Cuz you are great in explaining and coaching

OpenStudy (anonymous):

ahhh it is not true.....

OpenStudy (anonymous):

YOu gotta separate the t from any other constants. Make it t=

OpenStudy (anonymous):

4t=52

OpenStudy (anonymous):

rld m still 22 yrs old not so old dear.........

OpenStudy (anonymous):

Sheg don't blush from the compliment

OpenStudy (anonymous):

t+13=54

OpenStudy (anonymous):

no 4t = 52 now divide both sides by 4 and see what u get

OpenStudy (anonymous):

Bring the 4 to the other side of the equation. 4t=52 so divide t by 4 and 52 cus whatever u do to one side you gotta do to the other

OpenStudy (anonymous):

4t=52=13

OpenStudy (anonymous):

4t/4=t so t=52/4

OpenStudy (anonymous):

t=13

OpenStudy (anonymous):

great wlbrt u did it on ur own with sum guidence from rld......so, who did it u !!!!